# Parallel Batch-Dynamic Maximal Matching with Constant Work per Update

### v1

@@ Abstract.
We present a work optimal algorithm for parallel fully batch-dynamic maximal matching against an oblivious adversary. In particular it processes batches of updates (either insertions or deletions of edges) in constant expected amortized work per edge update, and in $O(\log^{3}m)$ depth per batch whp, where $m$ is the maximum number of edges in the graph over time. This greatly improves on the recent result by Ghaffari and Trygub (2024) that requires $O(\log^{9}m)$ amortized work per update and $O(\log^{4}m)$ depth per batch, both whp.
The algorithm can also be used for hyperedge maximal matching. For hypergraphs with rank $r$ (maximum cardinality of any edge) the algorithm supports batches of insertions and deletions with $O(r^{3})$ expected amortized work per edge update, and $O(\log^{3}m)$ depth per batch whp. This is a factor of $O(r)$ work off of the best sequential algorithm, Assadi and Solomon (2021), which uses $O(r^{2})$ work per update. Ghaffari and Trygub’s parallel batch-dynamic algorithm on hypergraphs requires $O(r^{8}\log^{9}m)$ amortized work per edge update whp. We leverage ideas from the prior algorithms but introduce substantial new ideas. Furthermore, our algorithm is relatively simple, perhaps even simpler than the sequential hyperedge algorithm.
We also present the first work-efficient algorithm for maximal matching on hypergraphs. For a hypergraph with total cardinality $m^{\prime}$ (i.e., sum over the cardinality of each edge), the algorithm runs in $O(m^{\prime})$ work in expectation and $O(\log^{2}m)$ depth whp. The algorithm also has some properties that allow us to use it as a subroutine in the dynamic algorithm to select random edges in the graph to add to the matching.
@@
ccs: Theory of computation Dynamic graph algorithms
@@ 1. Introduction
We consider the problem of maintaining the maximal matching of a graph under batch insertions and deletions of edges (fully dynamic), or hyperedges in a hypergraph. There have been several important results on sequential fully dynamic maximal matching in the past fifteen years. Onak and Rubinfeld described an dynamic algorithm with $O(\log^{2}n)$ expected amortized time per update that maintains a constant-approximate maximum matching (Onak and Rubinfeld, 2010). Baswana, Gupta and Sen (BGS) presented the first maximal matching algorithm with polylogarithmic time per update (Baswana et al., 2011). Their algorithm supports single updates (insertions or deletions) in $O(\log n)$ amortized time with high probability (whp). [^1] We say that $f(n)=O(g(n))$ with high probability in $n$ if for all constants $c>1$, there exist constants $n_{0},c_{0}>1$ such that for $n\geq n_{0}$, $Pr[f(n)\leq c_{0}g(n)]\geq 1-n^{-c}$. Solomon (Solomon, 2016) then improved the result to $O(1)$ time per update, again amortized and whp. These results are for standard graphs. Assadi and Solomon (AS) then extended the results to hypergraphs (Assadi and Solomon, 2021). In particular, they present an algorithm that, for a hypergraph where each edge has rank at most $r$, supports insertions and deletions in $O(r^{2})$ per update, amortized and whp. All of these results are against an oblivious adversary who knows the algorithm but must select the sequence of updates without knowing any of the random choices by the algorithm. All of these algorithms maintain a leveling structure that try to keep vertices or edges on a level appropriate for their degree.
@@
There exists a randomized algorithm that solves the parallel batch-dynamic maximal matching problem on hypergraphs with $O(r^{3})$ amortized and expected work per edge update. Furthermore, processing each batch requires $O(\log^{3}m)$ 0pt with high probability. Here $r$ is the rank of the hypergraph and $m$ is the maximum number of edges to appear at any point in the update sequence. The batch size can vary across steps.
@@
There exists a parallel randomized batch dynamic graph maximal matching algorithm with constant update time in expectation and polylog 0pt whp, which is work-optimal.
@@
We note that $r$-approximate set-cover for instances where each element belongs to at most $r$ sets reduces to maximal matching in a hypergraph with rank $r$ (Assadi and Solomon, 2021). In the reduction, sets correspond to vertices in the hypergraph and elements to the hyperedges. Our batch dynamic maximal matching algorithm therefore implies the following.
Corollary 1.3.
There exists a randomized parallel batch-dynamic $r$-approximate set cover algorithm that supports batches of insertions and deletions of elements, each element covered by at most $r$ sets. The algorithm does $O(r^{3})$ work per element update (expected and amortized) and has $O(\log^{3}m)$ 0pt per batch whp, where $m$ is the maximum number of elements to ever appear.
This is the current state of the art for the problem.
Within our algorithm, we use static hypergraph maximal matching as a subprocedure, and our work bound is proportional to the cost of static matching. However, directly translating the existing static parallel matching algorithms (Luby, 1986; Birn et al., 2013; Blelloch et al., 2012) to hypergraphs gives $O(mr^{2})$ work, or $O(mr\log m)$ work (Ghaffari and Trygub, 2024). Thus, to improve our work bounds, we develop the first work-efficient, and work optimal, parallel algorithm for maximal hypergraph matching. This builds on the work of Blelloch, Fineman and Shun (BFS) (Blelloch et al., 2012) and improved bounds by Fischer and Noever (FN)  (Fischer and Noever, 2019). FN show that the dependence 0pt of the simple greedy sequential algorithm for maximal matching is shallow when applied to a random permutation. This also applies to hypergraph matching. We extend the approach of BFS to develop a work-efficient algorithm. The cardinality of an edge in a hypergraph is the number of vertices it is incident on, and the total cardinality of a hypergraph is the sum of the cardinality across the edges.
@@ 1.1. Relationship to Prior Algorithms
Our algorithm borrows many ideas from prior level-based dynamic matching algorithms (Onak and Rubinfeld, 2010; Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021; Ghaffari and Trygub, 2024). It is most similar to the Assadi and Solomon (Assadi and Solomon, 2021) hypergraph algorithm. Like (Assadi and Solomon, 2021) and (Solomon, 2016), and unlike the others it is lazy: although a level accurately represents the original sample size when settled, the degree of an edge can grow much larger before it is resettled. This is what allows it to be work efficient.
@@
Our algorithm differs from prior algorithms in important ways, however, and adds several new ideas to achieve parallel work efficiency. One of the most important differences is the random sampling approach. Instead of selecting a single edge among the neighbors of a vertex (or in some special cases in (Assadi and Solomon, 2021), an edge), it selects the samples using a random greedy maximal matching algorithm, as described above. We show that the randomness in this very natural algorithm (the trivial greedy maximal matching on a random ordering of the edges) is sufficient.
@@ 2. Preliminaries
A hypergraph $H=(V,E)$ consists of a set of vertices $V$ and a set of edges $E\subset 2^{V}$. Let $n=|V|,m=|E|$ and $m^{\prime}=\sum_{e\in E}|e|$. The rank of a hypergraph is the maximum cardinality of any of its edges. We will be considering hypergraphs of bounded rank, and will use $r(H)$ or just $r$ to indicate the rank. We say two edges are incident on each other if they share a vertex, and an edge is incident on a vertex if it includes the vertex in its set. On a hypergraph $H=(V,E)$ a set of edges $M\subset E$ is said to be a matching if no two edges in $M$ are incident on each other. It is a maximal matching if all edges in $E\setminus M$ are incident on $M$. We say that an edge or vertex is free if it is not incident on any matched edge.
@@
We assume the fork-join parallel model with work-0pt analysis (Blelloch, 1996; Blelloch et al., 2020). In the model, a thread can fork off any number of child threads which run in parallel, but asynchronously. When all finish, control returns to the parent, which is the join point. Such forking can be nested. Beyond the fork instruction, each thread executes a standard RAM instruction set. We say there is a dependence between two instructions if they are in the same thread, or one is in an ancestor thread of the other, and otherwise they are concurrent. The cost is measured in terms of work and 0pt. The work is the total number of instructions executed by all threads, and the 0pt (also called span or critical path length) is the longest chain of dependent instructions. We allow for concurrent writes, but the ordering is non-deterministic. In particular, we assume that the values that are read are consistent with some total order of the instructions that is compatible with the partial order defined by the dependences.
A PRAM algorithm with $p$ processors and $t$ time can be simulated on the fork-join model in $O(pt)$ work and $O(t)$ 0pt. A fork-join algorithm with $w$ work and $t$ time can be simulated on the PRAM in $O(w/p+t\log^{*}w)$ time randomly w.h.p, or $O(w/p+t\log\log p)$ time deterministically (Blelloch et al., 2020). Because mapping from the fork-join model to the PRAM only adds at worst a $O(\log^{*}w)$ factor to the time in the randomized case, our results are quite robust across the models.
@@
We are interested in batch dynamic maximal matching. In particular we support inserting and deleting sets (batches) of edges. Each batch can be of arbitrary size. We also assume we can report for any vertex its matched edge (or none if the vertex is free) in constant expected time, and that we can return the set of matched edges in expected work proportional to the size of the set. We assume edges have unique identifiers so they can be hashed or compared for equality in constant time (even though they might have $r$ endpoints). This is not an unreasonable assumption since one can always keep a hash table that hashes an edge to its label external to our algorithm and then use the labels internally. As with much prior work on dynamic maximal matching (Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021; Ghaffari and Trygub, 2024) our algorithms are randomized and we assume an oblivious adversarial user. The oblivious adversary does not know anything about the random choices made by the algorithm.
@@
[…]
We also make use of dictionary data structures based on hashing. A parallel dictionary structure supports batches of insertions, deletions, and membership queries. A batch of any of these operations $k$ can be implemented in $O(k)$ expected work and $O(\log^{*}k)$ 0pt with whp (in $k$(Gil et al., 1991). This assumes the dictionary is large enough to hold the elements (i.e., the load factor is constant). We will need growable/shrinkable dictionaries, but given that we are happy with amortized bounds this can be implemented using the standard doubling/halving tricks (i.e., when too full, double the size and copy the old contents into the new, and when too empty, half the size and copy). In our analysis we will assume that batch of $k$ operations can be applied to a dictionary of size $n$ in $O(k)$ expected amortized work and $O(\log(n+k))$ 0pt whp. Also beyond membership queries we will need to extract a sequence of all the current elements. This can be performed with $O(n)$ expected work and $O(\log(n+k))$ 0pt whp. The size of each dictionary is $O(n)$.
Finally we will need a findNext function that, given an index $i$ in an array, finds the next index $j$ that satisfies some predicate. It needs to run in $O(j-i)$ work and $O(\log(j-i))$ 0pt. This can be implemented using doubling then binary search. On each round $k$, search the next $2^{k}$ elements. In a round, in parallel each element in the range checks the predicate and writes to some flag if satisfied. If none is found in a round, then go to the next round. Each round takes $O(2^{k})$ work and $O(1)$ 0pt. When a round succeeds, we then apply binary search over that range. In each step of binary search, we check if the left half of the list contains an alive element by concurrently writing to a flag. If the left half contains an alive element, we recurse on the left half; otherwise, we recurse on the right half. Thus the index of the next element that satisfies the predicate can be identified in $O(2^{k})$ work and $O(k)$ 0pt. The total work and 0pt is therefore within the required bounds.
@@
In our pseudocode, for simplicity, we often insert and delete elements from various sets in parallel loops. In our parallel model, this would require first applying a groupBy to gather elements based on the set they are updating, and then applying a batch update to each such set. Similarly, in the updateTop code in Section 3, we increment various counters in parallel. This can be implemented with a sumBy. Both of these require at most expected amortized work proportional to the number of updates, and logarithmic 0pt whp. The amortization comes from possibly needing to grow or shrink some of the hash tables for the sets.

### v2

@@ Abstract.
We present a work optimal algorithm for parallel fully batch-dynamic maximal matching against an oblivious adversary. It processes batches of updates (either insertions or deletions of edges) in constant expected amortized work per edge update, and in $O(\log^{3}m)$ depth per batch whp, where $m$ is the maximum number of edges in the graph over time. This greatly improves on the recent result by Ghaffari and Trygub (2024) that requires $O(\log^{9}m)$ amortized work per update and $O(\log^{4}m)$ depth per batch, both whp.
The algorithm can also be used for parallel batch-dynamic hyperedge maximal matching. For hypergraphs with rank $r$ (maximum cardinality of any edge) the algorithm supports batches of updates with $O(r^{3})$ expected amortized work per edge update, and $O(\log^{3}m)$ depth per batch whp. Ghaffari and Trygub’s parallel batch-dynamic algorithm on hypergraphs requires $O(r^{8}\log^{9}m)$ amortized work per edge update whp. We leverage ideas from the prior algorithms but introduce substantial new ideas. Furthermore, our algorithm is relatively simple, perhaps even simpler than Assadi and Solomon’s (2021) sequential dynamic hyperedge algorithm.
We also present the first work-efficient algorithm for parallel static maximal matching on hypergraphs. For a hypergraph with total cardinality $m^{\prime}$ (i.e., sum over the cardinality of each edge), the algorithm runs in $O(m^{\prime})$ work in expectation and $O(\log^{2}m)$ depth whp. The algorithm also has some properties that allow us to use it as a subroutine in the dynamic algorithm to select random edges in the graph to add to the matching.
With a standard reduction from set cover to hyperedge maximal matching, we give state of the art $r$-approximate static and batch-dynamic parallel set cover algorithms, where $r$ is the maximum frequency of any element, and batch-dynamic updates consist of adding or removing batches of elements.
@@
journalyear: 2025copyright: rightsretainedconference: 37th ACM Symposium on Parallelism in Algorithms and Architectures; July 28–August 1, 2025; Portland, OR, USAbooktitle: 37th ACM Symposium on Parallelism in Algorithms and Architectures (SPAA ’25), July 28–August 1, 2025, Portland, OR, USAdoi: 10.1145/3694906.3743313isbn: 979-8-4007-1258-6/25/07ccs: Theory of computation Dynamic graph algorithms
@@ 1. Introduction
We consider the problem of maintaining the maximal matching of a graph under batch insertions and deletions of edges (fully dynamic), or hyperedges in a hypergraph. There have been several important results on sequential fully dynamic maximal matching in the past fifteen years. Onak and Rubinfeld described an dynamic algorithm with $O(\log^{2}n)$ expected amortized time per update that maintains a constant-approximate maximum matching (Onak and Rubinfeld, 2010). Baswana, Gupta and Sen (BGS) presented the first maximal matching algorithm with polylogarithmic time per update (Baswana et al., 2011). Their algorithm supports single updates in $O(\log n)$ amortized time with high probability (whp). [^1] We say that $f(n)=O(g(n))$ with high probability in $n$ if there exists constants $n_{0},c$ such that for all constants $k>1$, we have for $n\geq n_{0}$, that $Pr[f(n)\leq ckg(n)]\geq 1-n^{-k}$. Solomon (Solomon, 2016) then improved the result to $O(1)$ time per update, again amortized and whp. These results are for standard graphs. Assadi and Solomon (AS) then extended the results to hypergraphs (Assadi and Solomon, 2021). First, they present an algorithm that, for a hypergraph where each edge has rank at most $r$, supports insertions and deletions in $O(r^{3})$ per update, amortized and whp. Then, they improve this algorithm to yield an $O(r^{2})$ per update algorithm, also amortized and whp. [^2] We believe there to be significant gaps in the $O(r^{2})$ algorithm, and are unsure of its correctness. We are in contact with the authors. All of these results are against an oblivious adversary who knows the algorithm but must select the sequence of updates without knowing any of the random choices by the algorithm. All of these algorithms maintain a leveling structure that try to keep vertices or edges on a level appropriate for their degree.
@@
There exists a randomized algorithm that solves the parallel batch-dynamic maximal matching problem on hypergraphs with $O(r^{3})$ amortized and expected work per edge update against an oblivious adversary. Furthermore, processing each batch requires $O(\log^{3}m)$ 0pt with high probability. Here $r$ is the rank of the hypergraph and $m$ is the maximum number of edges to appear at any point in the update sequence. The batch size can vary across steps.
@@
There exists a parallel randomized batch dynamic graph maximal matching algorithm with constant update time in expectation and polylog 0pt whp against an oblivious adversary, which is work-optimal.
@@
Cardinality matching has various applications (Burq, 2019; Wajc, 2020). Generally, the vertices of the graph model agents or resources, and edges connect compatible pairs. In a bipartite graph, one focuses on matching resources to consumers, in a general graph, one is pairing agents together, and in a hypergraph, one is selecting compatible groups of agents. Maximal matching accepts a smaller matching size in exchange for time savings. Batch-dynamic updates against an oblivious adversary allow the compatibility preferences of agents to change over time, but only due to outside effects (in ways agnostic to the matching itself). In summary, our setting, parallel batch-dynamic hypergraph maximal matching against an oblivious adversary, addresses the problem of selecting groups of agents quickly, with compatibility preferences that change over time only due to outside effects, where speed is more important than optimizing the number of groups matched. Theoretically, the attention (Onak and Rubinfeld, 2010; Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021) the sequential dynamic algorithms community has given to maximal matching against an oblivious adversary renders it interesting in the parallel-batch dynamic setting. In turn, matching received attention from the dynamic algorithms community perhaps because of static matching’s relevance in TCS as a whole; for example, Edmonds’s Blossom algorithm (for finding a maximum matching on a general graph) contributed to the use of polynomial time as a measure of efficiency (Edmonds, 1965; Mehta, 2013).
Within our algorithm, we use static hypergraph maximal matching as a subprocedure, and our work bound is proportional to the cost of static matching. However, directly translating the existing static parallel matching algorithms (Luby, 1986; Birn et al., 2013; Blelloch et al., 2012) to hypergraphs gives $O(mr^{2})$ work, or $O(mr\log m)$ work (Ghaffari and Trygub, 2024). Thus, to improve our work bounds, we develop the first work-efficient, and work optimal, parallel algorithm for maximal hypergraph matching. This builds on the work of Blelloch, Fineman and Shun (BFS) (Blelloch et al., 2012) and improved bounds by Fischer and Noever (FN)  (Fischer and Noever, 2019). These works show that the dependence 0pt of the simple greedy sequential algorithm for maximal matching is shallow when applied to a random permutation. This also applies to hypergraph matching. We extend the approach of BFS to develop a work-efficient algorithm. The cardinality of an edge in a hypergraph is the number of vertices it is incident on, and the total cardinality of a hypergraph is the sum of the cardinalities of the edges.
@@
We note that $r$-approximate set-cover for instances where each element belongs to at most $r$ sets reduces to maximal matching in a hypergraph with rank $r$ (Assadi and Solomon, 2021). In the reduction, sets correspond to vertices in the hypergraph and elements to the hyperedges. Our static and batch dynamic maximal matching algorithms therefore imply the following corollaries.
Corollary 1.4.
There exists a randomized parallel batch-dynamic $r$-approximate set cover algorithm that supports batches of insertions and deletions of elements against an oblivious adversary, each element covered by at most $r$ sets. The algorithm does $O(r^{3})$ work per element update (expected and amortized) and has $O(\log^{3}m)$ 0pt per batch whp, where $m$ is the maximum number of elements to ever appear.
Corollary 1.5.
There exists a randomized algorithm for static parallel $r$-approximate set cover, with $O(m^{\prime})$ expected work and $O(\log^{2}m)$ depth whp, where $m^{\prime}$ is the total cardinality of the $m$ edges in the graph, and where no element appears in more than $r$ sets.
These algorithms are the current parallel static and parallel batch-dynamic state of the art for $r$-approximate set cover. Note that there are two types of set cover approximations, those based on the maximum frequency of any element ($r$), and those based on the maximum size of any set ($\Delta$) (Dhulipala et al., 2024). There are many parallel results with good work bounds, but which sacrifice a $(1+\epsilon)$ factor in the approximation ratio; hypergraph matching is unusual in providing precisely an $r$-approximate cover. Thus, our set cover bounds improve on the bounds given by previous parallel matching algorithms (Blelloch et al., 2012; Ghaffari and Trygub, 2024). For more information on parallel set cover, see Dhulipala et al. (Dhulipala et al., 2024).
For the full version of the paper, see arXiv (Blelloch and Brady, 2025).
@@ 1.1. Relationship to Prior Algorithms
Our algorithm borrows many ideas from prior level-based dynamic matching algorithms (Onak and Rubinfeld, 2010; Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021; Ghaffari and Trygub, 2024). It is most similar to the Assadi and Solomon (Assadi and Solomon, 2021) hypergraph algorithm. Like (Assadi and Solomon, 2021) and (Solomon, 2016), and unlike the others it is lazy: although a level accurately represents the original sample size of an edge when settled, the degree of an edge can grow much larger before it is resettled. This is what allows it to be work efficient.
@@
As in previous parallel batch-dynamic algorithms, we use a parallel static maximal matching on insertions to determine which edges to match (Liu et al., 2022; Ghaffari and Trygub, 2024). Our algorithm differs from prior algorithms in important ways, however, and adds several new ideas to achieve parallel work efficiency. One of the most important differences is the random sampling approach. Instead of selecting a single edge among the neighbors of a vertex (or in some special cases in (Assadi and Solomon, 2021), an edge), it selects the samples using a random greedy maximal matching algorithm across many vertices and edges. We show that the randomness in this very natural algorithm (the trivial greedy maximal matching on a random ordering of the edges) is sufficient.
@@ 2. Preliminaries
A hypergraph $H=(V,E)$ consists of a set of vertices $V$ and a set of edges $E\subset 2^{V}$. Let $n=|V|,m=|E|$ and $m^{\prime}=\sum_{e\in E}|e|$. The rank of a hypergraph is the maximum cardinality of any of its edges. We will be considering hypergraphs of bounded rank, and will use $r(H)$ or just $r$ to indicate the rank. We say two edges are incident on each other if they share a vertex, and an edge is incident on a vertex if it includes the vertex in its set. Two edges are neighbors if they are incident on each other. On a hypergraph $H=(V,E)$ a set of edges $M\subset E$ is said to be a matching if no two edges in $M$ are incident on each other. It is a maximal matching if all edges in $E\setminus M$ are incident on $M$. We call an edge or vertex free if it is not incident on any matched edge.
@@
We assume the binary forking model with work-0pt analysis (Blelloch, 1996; Blelloch et al., 2020). In the model, a thread can fork off any number of child threads which run in parallel, but asynchronously. When all finish, control returns to the parent, which is the join point. Such forking can be nested. Beyond the fork instruction, each thread executes a standard RAM instruction set. We say there is a dependence between two instructions if they are in the same thread, or one is in an ancestor thread of the other, and otherwise they are concurrent. The cost is measured in terms of work and 0pt. The work is the total number of instructions executed by all threads, and the 0pt is the longest chain of dependent instructions. We allow for concurrent writes (with constant work per write, and constant 0pt), but the ordering is non-deterministic. In particular, we assume that the values that are read are consistent with some total order of the instructions that is compatible with the partial order defined by the dependences. A PRAM algorithm with $p$ processors and $t$ time can be simulated on the binary forking model in $O(pt)$ work and $O(t)$ 0pt. A binary forking algorithm with $w$ work and $t$ time can be simulated on the PRAM in $O(w/p+t\log^{*}w)$ time randomly w.h.p, or $O(w/p+t\log\log p)$ time deterministically (Blelloch et al., 2020). Because mapping from the binary forking model to the PRAM only adds at worst a $O(\log^{*}w)$ factor to the time in the randomized case, our results are quite robust across the models.
@@
We are interested in batch-dynamic maximal matching. In particular we support inserting and deleting sets (batches) of edges. Each batch can be of arbitrary size. We also assume we can report for any vertex its matched edge (or none if the vertex is free) in constant expected time, and that we can return the set of matched edges in expected work proportional to the size of the set. We assume edges have unique identifiers so they can be hashed or compared for equality in constant time (even though they might have $r$ endpoints). This is not an unreasonable assumption since one can always keep a hash table that hashes an edge to its label external to our algorithm and then use the labels internally. As with much prior work on dynamic maximal matching (Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021; Ghaffari and Trygub, 2024) our algorithms are randomized and we assume an oblivious adversarial user. The oblivious adversary does not know anything about the random choices made by the algorithm.
@@
We will make significant use of semisorting (Valiant, 1990). The semisorting problem is given a set of keys, to organize them so equal keys are adjacent in the output, but not necessarily sorted. For semisorting we assume we have a hash function on the keys. Semisorting a collection of $n$ keys requires $O(n)$ expected work and $O(\log n)$ 0pt whp  (Valiant, 1990; Gu et al., 2015). Semisort can, in turn, be used for several other operations. In particular we will need a groupBy function that takes a set of key-value pairs and groups all values with the same key into its own subset. It returns a set of pairs, each consisting of a unique key and all the values with that key. This is easily implemented by a semisort followed by a prefix sum to partition into subsets. We will also need a sumBy operation that takes a set of key-value pairs where the values are integers, and for each unique key sums up the values with that key. It returns a set of key-value pairs where the keys are unique and the values are the corresponding sums. This again can be implemented with a semisort. We also need a removeDuplicates function, that given a list with repeated elements, returns the unique list. This can be implemented with a groupBy. In our pseudocode, when we use set notation $\{\}$, we are implicitly calling removeDuplicates as needed.
We also make use of dictionary data structures based on hashing. A parallel dictionary structure supports batches of insertions, deletions, and membership queries. A batch of any of these operations $k$ can be implemented in $O(k)$ expected work and $O(\log^{*}k)$ 0pt whp in $k$ (which is $O(\log n)$ 0pt whp in $n$)  (Gil et al., 1991). This assumes the dictionary is large enough to hold the elements (i.e., the load factor is constant). We will need growable/shrinkable dictionaries, but given that we are happy with amortized bounds this can be implemented using the standard doubling/halving tricks (i.e., when too full, double the size and copy the old contents into the new, and when too empty, half the size and copy). In our analysis we will assume that batch of $k$ operations can be applied to a dictionary of size $n$ in $O(k)$ expected amortized work and $O(\log(n+k))$ 0pt whp. Also beyond membership queries we will need to extract a sequence of all the current elements. This can be performed with $O(n)$ expected work and $O(\log(n+k))$ 0pt whp. The size of each dictionary is $O(n)$.
Finally we will need a findNext function that, given an index $i$ in an array, finds the next index $j$ that satisfies some predicate. It needs to run in O(ji)O(j-i) work and $O(\log(j-i))$ 0pt. This can be implemented using doubling then binary search. On each round $k$, search the next $2^{k}$ elements. In a round, in parallel each element in the range checks the predicate and writes to a flag if satisfied. [^4] We can do this because we support concurrent writes with a non-deterministic ordering. If none is found in a round, then go to the next round. Each round takes $O(2^{k})$ work and $O(1)$ 0pt. When a round succeeds, we then apply binary search over that range. In each step of binary search, we check if the left half of the list contains an alive element by concurrently writing to a flag. If the left half contains an alive element, we recurse on the left half; otherwise, we recurse on the right half. Thus the index of the next element that satisfies the predicate can be identified in $O(2^{k})$ work and $O(k)$ 0pt. The total work and 0pt is therefore within the required bounds.
@@
In our pseudocode, for simplicity, we often insert and delete elements from various sets and maps in parallel loops where the parallel iterations might be on the same key. In our parallel model, this would require first applying a groupBy to gather elements based on the set they are updating, and then applying a batch update to each such set. Similarly, in the updateTop code in Section 3, we increment various counters in parallel. This can be implemented with a sumBy. Both of these require at most expected amortized work proportional to the number of updates, and logarithmic 0pt whp. The amortization comes from possibly needing to grow or shrink some of the hash tables for the sets.